OpenStudy (anonymous):

Verify the identity, show all work tanθ cotθ-sin^2 θ=cos^2 θ

OpenStudy (psymon):

tanθ = sinθ/cosθ and cotθ = cosθ/sinθ. So I'll rewrite your problem now (sin/cos)*(cos/sin) - sin^2θ = cos^2θ The multiplication cancels and becomes 1. So now you have your pythagorean theorem identity, 1- sin^2θ = cos^2θ, which is of course true.

OpenStudy (anonymous):

@Psymon can you help me with these 3? All that you can/ want to do anyway 1/(1+sinx)+1/1+cscx)=1 (1-2sin^2 θ)/(sinθcosθ)=cotθ-tanθ √([1+sinθ]/[1-sinθ])=(1+sinθ)/|cosθ| thank you so, so much!

OpenStudy (anonymous):

^^those are all for verifying the identity